A function is given by the equation 𝑦 = 2π‘₯ 3 + 4π‘₯ 2 βˆ’ 5π‘₯ + 10.

i) Obtain an expression for 𝑑𝑦 𝑑π‘₯.

ii) Find the gradient of the tangent to the curve at (2, 6).

iii) Determine the equation of this tangent line which passes through (2, 6)


1
Expert's answer
2021-11-21T17:13:15-0500

"y=2x\u00b3+4x\u00b2-5x+10"



(i) "\frac{dy}{dx}=6x²+8x-5"



(ii) gradient at (2, 6)


"=> x=2, y = 6"


"\frac{dy}{dx}=6x²+8x-5=6(2)²+8(2)-5=35"



(iii) "\frac{dy}{dx}=m=\frac{y-y_1}{x-x_1}"


"=> 35=\frac{y-6}{x-2}"


"=> 35(x-2)=(y-6)"


"=>35x-70=y-6"


"=> y-35x+64=0"